### v1

@@ 3 Hessian-free Laplace (HFL)
$\nabla_{\lambda}f_{\hat{\theta}^{(x,\lambda)}}(x)\biggr{\rvert}_{\lambda=0}=\nabla_{\theta}f_{\hat{\theta}}(x)^{\top}P^{-1}\nabla_{\theta}f_{\hat{\theta}}(x),$ (14)
assuming $\hat{\theta}^{(x,\lambda)}$ and $P^{-1}$ exist, and that ${\mathcal{L}_{\theta}}^{(x,\lambda)}$ is continuously differentiable w.r.t. $\theta$.
@@
Consider the derivative w.r.t. $\theta$ of the prediction-regularized objective at the optimum,
@@
A parallel frequentist approach is known as the implicit delta method Kallus & McInerney (2022). The relationship in Eq. 14 can also be understood as an application of the implicit function theorem (IFT) by Cauchy, e.g. see Lorraine et al. (2020) for a modern treatment of IFT.
$\frac{1}{\lambda}(f_{\hat{\theta}^{(x,\lambda)}}(x)-f_{\hat{\theta}}(x))\underset{\lambda\to 0}{\rightarrow}\nabla_{\lambda}f_{\hat{\theta}^{(x,\lambda)}}(x)\biggr{\rvert}_{\lambda=0}.$ (19)
It is natural to consider auto-differentiation to calculate the LHS of Eq. 14. From the fact that the optimum $\hat{\theta}^{(x,\lambda)}$ also changes with $\lambda$ it can be seen that auto-differentiation must deal with complex operations on optima and is unlikely to lead to a computational advantage unless used in tandem with additional work-arounds.
Putting it together, the Hessian-free Laplace (HFL) approximation is,
$\tilde{f}_{\theta}\sim\mathcal{N}\left(f_{\hat{\theta}}(x),\frac{1}{\lambda}(f_{\hat{\theta}^{(x,\lambda)}}(x)-f_{\hat{\theta}}(x))\right),$ (20)
and differs from linearized Laplace (Eq. 11) by the lack of an explicit Hessian. It is apparent by this comparison that HFL trades the computational and storage expense of dealing with the Hessian for that of a point-wise approximation for each query $x$. There are strong arguments in favor of the latter. First, for suitably small $\lambda$, the prediction-regularized MAP is close to the MAP for any choice of $x$, and so relatively inexpensive to obtain with stochastic gradient descent. By comparison, matrix inversion takes $\mathcal{O}(K^{3})$ computations in the number of network parameters $K$. Second, HFL sidesteps the issue of which Hessian approximation to use, a problem analogous to hyperparameter selection in that it depends on several factors including network architecture, likelihood function, and dataset. The approach is summarized in Algorithm 1 and uses the stochastic objective,
~$\tilde{\mathcal{L}}_{\theta}^{(x,\lambda)}(x_{i},y_{i})=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{n}f_{\theta}(x),$ (21)
for training example $(x_{i},y_{i})$, or an average over a minibatch.
3.1 Pre-Training Hessian-Free Laplace
Algorithm 2 Pre-Training Hessian-Free Laplace
  Input: network $f_{\theta}$, fitted MAP $\hat{\theta}$, training data $\mathcal{D}$ of size $n$, evaluation inputs $\{\hat{x}_{m=1}^{M}\}$, scalar $\lambda$, step size schedule $\gamma(\cdot)$
  Warm start $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}$
  Initialize step counter $j=0$
  repeat
     Sample training example $(x_{i},y_{i})\sim\mathcal{D}$
     Follow gradient $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}^{(\hat{r},\lambda)}+n\gamma(j)\nabla_{\theta}\mathcal{L}_{i}\;\;\;$ //Choose $\mathcal{L}_{i}$ from Eq. 25272829, or 31.
     $j\leftarrow j+1$
  until $\hat{\theta}^{(\hat{r},\lambda)}$ converges
  Return: predictive variance function $\hat{\sigma}_{f}^{2}(x):=\frac{1}{\lambda}|f_{\hat{\theta}^{(\hat{r},\lambda)}}(x)-f_{\hat{\theta}}(x)|$ (or parameter variance $\hat{\sigma}_{\theta}^{2}(x):=\frac{1}{\lambda}|\hat{\theta}^{(\hat{r},\lambda)}-\hat{\theta}|$)
HFL targets the same predictive variance-covariance as the Laplace approximation. However, when evaluating the uncertainty of many predictions, the requirement to fine-tune to the $f$-objective specific to each test point is a bottleneck. In order to scale HFL in the number of test points, we develop a pre-training approach that requires only a single model to represent the $f$-regularized parameters for a set of evaluation points.
Pre-training is formulated as follows. Find the regularization term $\hat{r}$ for a finite set of evaluation inputs ${\bf\hat{x}}=\{\hat{x}_{1:M}\}$ that minimizes the average squared L2 norm error of the corresponding $f$-regularized parameters,
$\hat{r}=\arg_{r}\min\sum_{m=1}^{M}\|\hat{\theta}^{(f(x_{m}),\lambda)}-\hat{\theta}^{(r,\lambda)}\|_{2}^{2},$ (22)
using expanded notation for $\hat{\theta}^{(\cdot,\lambda)}$ to accommodate different functions in the regularization term. Eq. 22 facilitates the calculation of Eq. 20 using $\hat{\theta}^{(\hat{r},\lambda)}$ instead of $\hat{\theta}^{(f(x),\lambda)}$ for arbitrary $x$ in the domain of $f$.
Theorem 3.2.
The optimal regularization term defined by Eq. 22 is $\hat{r}=\frac{1}{M}\sum_{m=1}^{M}f(\hat{x}_{m})$ when $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist and ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
@@ 3.1 Pre-Training Hessian-Free Laplace
A stochastic optimization algorithm for pre-training HFL is given in Algorithm 2, with objective,
$\mathcal{L}_{i}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nM}\sum_{m=1}^{M}f_{\theta}(\hat{x}_{m}).$ (25)
This conceptually simple procedure avoids the need to explicitly calculate the Hessian by optimizing one parameter vector $\hat{\theta}^{(\hat{r},\lambda)}$. The evaluation points may be the training dataset, or a specific test set which may be under dataset shift or require out-of-distribution (OOD) prediction.
There are four noteworthy variants based on this pre-training approach that we discuss next. Each can be used in the gradient update step of Algorithm 2, though their applicability varies, as we will discuss.
Absolute-Valued Regularizer
HFL can be reformulated to absorb a sign term into $\lambda$ that depends on the sign of the output function by defining,
$\tilde{\lambda}=\lambda(-1)^{\mathbb{I}[f_{\theta}(x)<0]},$ (26)
where $\mathbb{I}[c]$ is the indicator function that evaluates to 1 when the condition $c$ is true and to 0 otherwise. Equivalently, we can use the absolute values of both the HFL regularizer and predictive variance. This leads to the pre-training objective,
$\mathcal{L}_{i}^{\mathrm{(AV)}}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nM}\|f_{\theta}({\bf\hat{x}})\|_{1},$ (27)
where the final term in Eq. 27 is the (rescaled) sum of absolute values of the function applied to the $M$ evaluation inputs.
In-Sample Pre-Training & Data Augmentation
When the train and test data distributions are the same, the minibatches sampled to optimize the log joint probability of the model may also be used to regularize it, giving rise to the single-sample stochastic gradient,
$\mathcal{L}_{i}^{\mathrm{(IS)}}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{\lambda}{n}|f_{\theta}(x_{i})|+\frac{1}{nM}\log p(\theta).$ (28)
Furthermore, when the likelihood is Gaussian, the extra term can be absorbed into the likelihood. In particular, completing the square in Eq. 28, by pulling the term $\frac{\lambda\sigma^{2}}{nM}(\frac{\lambda\sigma^{2}}{2nM}-y_{i})$ from a constant that does not depend on $\theta$, yields the objective,

### v2

@@ 4.1 Amortized Regularizer
The optimal regularization term defined by Eq. 13 is $\hat{r}=\frac{1}{m}\sum_{i=1}^{m}f(\hat{x}_{i})$ when $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist and ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
@@
See Appendix A. ∎
Theorem 4.2 implies that embedding the average network output in the training objective minimizes the L2 norm of the $f$-regularized objective. Furthermore, we can absorb a sign term into $\lambda$ that depends on the sign of the output function by defining,
$\tilde{\lambda}=\lambda(-1)^{\mathbb{I}[f_{\theta}(x)<0]},$ (14)
where $\mathbb{I}[c]$ is the indicator function that evaluates to 1 when the condition $c$ is true and to 0 otherwise. This equates to using the absolute values of both the RegVar regularizer and predictive variance. Putting it together leads to the amortized objective,
$\mathcal{L}_{i}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nm}\|f_{\theta}({\bf\hat{x}})\|_{1},$ (15)
where the final term in Eq. 15 is the (rescaled) sum of absolute values of the function applied to the $m$ evaluation inputs. We emphasize that this sum of absolute values accommodates networks with multiple outputs, a crucial requirement for practical use.
A stochastic optimization algorithm for amortized RegVar is given in Algorithm 1. The procedure avoids the need to explicitly calculate the Hessian and instead optimizes one parameter vector $\hat{\theta}^{(\hat{r},\lambda)}$. The general form of the objective optimized in the algorithm is Eq. 15. We next highlight a special case and a variant of Eq. 15, both of which admit simplified forms of RegVar.
Special Case: In-Sample Amortization
When the train and test data distributions are the same, the data sampled to optimize the log joint probability of the model may also be used to regularize it, giving rise to the single-sample stochastic gradient,
$\mathcal{L}_{i}^{\mathrm{(IS)}}$ $=\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\frac{1}{n}\log p(\theta)+\frac{\lambda}{nm}|f_{\theta}(x_{i})|.$ (16)
Alternatively, a data augmentation approach is applicable to in-sample training when the likelihood is Gaussian (see Appendix B).
Variant: Parameter Uncertainty Quantification
The same amortized regularization strategy also enables parameter uncertainty quantification by finding the regularizer $\hat{t}$ that minimizes the average squared L2 norm over the $K$ dimensions of parameters,
Theorem 4.2 applies to Eq. 17 in combination with absolute value regularization to provide the amortized objective,
which corresponds to a set of sparsity-inducing Laplace priors over the weights. Thus, when used in concert with a Laplace log prior $\log p(\theta)$, the regularization term for RegVar is absorbed, with $\lambda$ perturbing the precision parameter. In this case, it is particularly convenient to implement RegVar for parameter uncertainty, as one needs only to adjust the strength of the existing regularizer, as detailed in Appendix B.3.
Table 1: Held-Out Predictive Errors with GPT-3, XL GPT-3, and ViT Architectures
Method GPT-3 (125M Param.) XL GPT-3 (1.3B Param.) ViT (86M Param.)
Wikitext IMDB Wikitext IMDB Places365
Negative Log Likelihood ($\downarrow$ is better) ±$\pm 95\%$ Confidence Interval
Pre-Trained $9.781\pm 0.009$ $10.374\pm 0.012$ $8.228\pm 0.013$ $9.886\pm 0.015$ $5.900\pm 0.002$
MAP $1.383\pm 0.027$ $9.681\pm 0.015$ $1.523\pm 0.030$ $9.206\pm 0.020$ $5.871\pm 0.012$
Ensemble $3.925\pm 0.058$ $22.84\pm 0.005$ $5.885\pm 0.010$
MC Dropout $1.410\pm 0.026$ $9.551\pm 0.013$ $1.495\pm 0.029$ $9.137\pm 0.019$ $5.878\pm 0.006$
Laplace Diag $1.481\pm 0.021$ $9.090\pm 0.011$ $1.273\pm 0.021$ $8.797\pm 0.017$ $5.872\pm 0.026$
Laplace Full $1.277\pm 0.021$ $8.774\pm 0.009$ $1.198\pm 0.021$ ${8.047}\pm 0.013$ $5.869\pm 0.001$
RegVar (this paper) ${1.143}\pm 0.021$ ${8.009}\pm 0.008$ ${1.151}\pm 0.020$ $8.138\pm 0.007$ ${5.853}\pm 0.014$
Expected Calibration Error ($\downarrow$ is better) ±$\pm 95\%$ Confidence Interval
Pre-Trained $0.163\pm 0.001$ $0.239\pm 0.001$ $0.233\pm 0.002$ $0.287\pm 0.002$ $0.001\pm 0.000$
MAP $0.052\pm 0.001$ $0.224\pm 0.002$ $0.060\pm 0.001$ $0.240\pm 0.002$ $0.009\pm 0.001$
Ensemble $0.153\pm 0.002$ $0.747\pm 0.004$ $0.002\pm 0.000$
MC Dropout $0.036\pm 0.000$ $0.176\pm 0.002$ $0.054\pm 0.001$ $0.219\pm 0.002$ $0.005\pm 0.000$
Laplace Diag $0.180\pm 0.002$ $0.170\pm 0.001$ $0.072\pm 0.001$ $0.203\pm 0.001$ $0.001\pm 0.000$
Laplace Full $0.082\pm 0.001$ $0.136\pm 0.001$ ${0.024\pm 0.000}$ $0.126\pm 0.001$ ${0.000\pm 0.000}$
RegVar (this paper) ${0.019\pm 0.000}$ ${0.044\pm 0.001}$ $0.047\pm 0.001$ ${0.024\pm 0.001}$ $0.001\pm 0.000$
Figure 3: Calibration curves for the methods on the IMDB held-out evaluation with the XL GPT-3 network. Ideal calibration is plotted with dashed lines. 95% confidence intervals are shown with horizontal markers.
5 EXPERIMENTS
We investigate the empirical performance of RegVar in quantifying uncertainty for large models. The goal is to validate the theoretical developments from Section 3 and compare to baselines in Bayesian deep learning. We next describe the experimental setup.
Experimental Setup
The experiments use a Small GPT-3 Architecture language model with 125 million parameters and a Vision Transformer (Base) with 86 million parameters. To study how the methods scale, we also experiment with a larger XL GPT-3 Architecture comprising 1.3 billion parameters. For the language models, we load the pre-trained model and fine-tune it to Wikitext (wikitext-103-raw) comprising unprocessed text extracted from Wikipedia (Merity et al., 2016) and IMDB a binary sentiment classification dataset comprising movie reviews split evenly between positive and negative sentiments (Maas et al., 2011). For the vision model, we fine-tune the pre-trained model to Places365 (places365), an image dataset of 365 scenes (Zhou et al., 2017). The fine-tuning language datasets were subsampled to 40k documents split into 80% train, 10% validation, and 10% test, while the vision dataset was subsampled to 4k images split into 50% train, 25% validation, and 25% test.
The language models were fine-tuned over 1 epoch with stochastic gradient descent using adaptive learning rates, [^1] Adam with learning rate $2\times 10^{-5}$ (Kingma and Ba, 2017). weight decay $0.01$, [^2] Equivalent to $\mathcal{N}(0,50I)$ prior on network parameters. and batch size 4 with 4 gradient accumulation steps. These experiments were conducted on an instance configured with one NVIDIA A100 GPU with 40 GB memory, 12 CPU cores, and 143 GB of system memory. The vision model was fine-tuned over 5 epochs with the same settings without the gradient accumulation on an instance configured with 4 NVIDIA A10 GPUs each with 24 GB memory, 96 CPU cores, and 380 GB of system memory.
Figure 4: Contour plot of held-out negative log likelihood ($\downarrow$ is better) for Wikitext validation as function of the $\lambda$ used to train RegVar and variance rescaling used during inference. Lines isometric to the identity line are shown as gray dashes.
We compared a number of uncertainty quantification methods,
  • Maximum a posteriori (MAP) optimizes the log joint distribution w.r.t. the network parameters (see Eq. 1).
  • Ensemble Last-Layer performs MAP estimation with multiple heads at the last layer with a shared network before the last layer (Lakshminarayanan et al., 2017; Osband and Van Roy, 2015). During training, heads are randomly initialized and, for each training example, a head is sampled at random to receive an update. During inference, the mean and variance of the population of all head evaluations is used for uncertainty. Memory constraints prohibit approximations beyond the last layer and prevent applying ensembling to the XL GPT-3 architecture.
  • Monte Carlo Dropout (MC Dropout) uses dropout to generate approximate samples of the network during inference time (Gal and Ghahramani, 2016). Since dropout is applied at every layer, it is the only other method (apart from RegVar) that considers uncertainty beyond the last layer.
  • Last-Layer Laplace with Diagonal Covariance (Laplace Diag.) using the inverse of the diagonal of the GGN approximation of the Hessian of the parameters in the last layer of the network.
  • Last-Layer Laplace with Full Covariance (Laplace Full) using the inverse of the GGN approximation of the Hessian of the parameters in the last layer of the network.
  • Regularization Variation (RegVar) uses Algorithm 1.
Further variants of the Laplace approximation can be brought to bear on large models (Daxberger et al., 2021). Last-layer diagonal/full covariance Laplace are used here because we found them, in line with previous studies, to be representative of the performance of different Laplace approaches.
Hyperparameters
All network weights are regularized by a Gaussian prior with zero mean and low precision (0.02) during fine-tuning. Due to computational constraints, 10 samples are used in the ensemble and MC dropout methods. Dropout rate, prior conditioning matrix for Laplace, and the $\lambda$ weight in RegVar are tuned to maximize likelihood on validation data. For all methods there is a trade-off between accuracy and calibration; to control for the confounding influence of this trade-off on the metrics, we rescale the predicted variance by a scalar hyperparameter that is chosen for each method to maximize likelihood on the validation datasets.
Evaluation
All methods are evaluated by held-out negative log likelihood (NLL) and expected calibration error (ECE). NLL penalizes predictive uncertainties that place more probability on the wrong classes, indicating both accuracy and calibration. ECE measures calibration directly using the weighted average error of binned predictions,
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|,$ (19)
where $M$ is the total number of bins, $B_{m}$ is the set of samples whose predicted confidence scores fall into the $m$-th bin, $n$ is the total number of samples, $\text{acc}(B_{m})$ is the accuracy of the samples in the $m$-th bin, and $\text{conf}(B_{m})$ is the average predicted confidence of the samples in the $m$-th bin. We use $M=10$ throughout. Confidence intervals are calculated using ±$\pm 1.96$ standard errors (Wald confidence intervals).
For every method, the first step to evaluating the predictive distribution over classes is to estimate the logit means $f_{\theta}(x_{i})$ and variances $\hat{\sigma}^{2}_{i}$ (rescaled by the aforementioned hyperparameter). Uncertainty in the logits is then used in the extended probit approximation (Gibbs, 1998) which introduces a logit correction vector $\kappa$ to account for model variance. The probit approximation correction is given by,
$\kappa_{i}=\frac{1}{\sqrt{1+\frac{\pi\hat{\sigma}^{2}_{i}}{8}}}.$ (20)
The adjusted logits are then used to form predicted probabilities over the clases $p(y_{i}\;|\;\mathcal{D},x_{i})\propto\exp\{\kappa_{i}f_{\theta}(x_{i})\}$.
Figure 5: Fine-tune training and inference times for the methods on IMDB dataset. N.B. ensemble was prohibited in the larger model due to memory constraints.
The implementation was done in PyTorch (Paszke et al., 2017) using HuggingFace333huggingface.co to load pre-trained models from EleutherAI/{gpt-neo-125M,gpt-neo-1.3B} repositories (Black et al., 2021) originally trained on a 800GB text dataset (Gao et al., 2020) and google/vit-base-patch16-224-in21k repositories (Wu et al., 2020) originally trained on ImageNet-21k (Deng et al., 2009). The code to run the experiments can be accessed publicly on GitHub. [^4] https://github.com/jamesmcinerney/regvar
Results
A comparison of evaluation metrics for the methods is shown in Table 3. We find that RegVar offers significant advantages over MAP across datasets and improves or maintains uncertainty quantification in comparison to the other methods. While Laplace Full (Last Layer) performs robustly across settings, improvements with RegVar highlight the potential benefits of deep uncertainty, i.e., including the sources of variance from the entire network, which is not tractable to target with explicit Hessians at this scale. While NLL measures a combination of accuracy and calibration, ECE isolates the contribution of calibration and allows us to explore how this breaks down further with calibration curves in Figure 3. Overall calibration is good, as evidenced by the small deviation from the ideal line, however most methods tend to over-represent extreme predictions (at 0/1) while RegVar is able to calibrate the predictions away from the extremes.
We explore the empirical behavior of the $\lambda$ hyperparameter in RegVar by running fine-tuning with a wide range of different values. Figure 4 shows a contour plot of the effect of $\lambda$ (x-axis) and the variance rescaling hyperparameter (y-axis), the latter a hyperparameter used in all methods. The isometry of the contours suggests insensitivity of held-out negative log likelihood to the choice of $\lambda$ during fine-tuning. The two factors appear to cancel each other out and leave a dataset-dependent constant rescaling term (approx. $\frac{1}{100}$ for Wikitext) that does not depend on $\lambda$.
A wall clock time comparison for fine-tuning and inference is shown in Figure 5. While inference performance of RegVar is advantageous to all other methods apart from MAP, obtaining the regularized network in RegVar doubles the MAP fine-tuning time, a cost that is amenable to further optimization (e.g. warm starting from MAP) and parallelization in a similar fashion to ensembling and dropout.
Additional experiments are detailed in Appendix C, exploring further settings including using the RegVar framework for parameter uncertainty.
Limitations
RegVar is inherently a local method centered around the optimum of the joint distribution and the curvature is evaluated at one location (ignoring other modes). Early stopping during fine-tuning or pre-training also means that the chosen network parameters may not be around the mode of the explicit loss. The finite differences required by RegVar are an additional source of error making it important to select $\lambda$ small enough, but not so small that underflow errors arise or, more practically, that the noise in the gradients overwhelm the signal from the regularizer.
6 CONCLUSIONS & FUTURE WORK
In this paper, we explored a new concept in Bayesian deep learning that identifies variation due to regularization as an approximation to the predictive variance. We considered this phenomenon from a linearized Laplace perspective and evaluated it in large models. There is further work to explore in what cases it is necessary to use the uncertainty of the whole network for predictions to help contextualize the value of RegVar.
ACKNOWLEDGEMENTS
We thank the anonymous AISTATS reviewers for their insightful feedback and suggestions.
Appendix A PROOFS
We detail here the proofs of the theorems given in the main text. Table 2 summarizes the notation used.
Theorem 4.1 The derivative of the prediction under the prediction-regularized MAP w.r.t. $\lambda$ recovers the variance-covariance term in Eq. 7,
$\nabla_{\lambda}f_{\hat{\theta}^{(f(x),\lambda)}}(x)\biggr{\rvert}_{\lambda=0}=\nabla_{\theta}f_{\hat{\theta}}(x)^{\top}P^{-1}\nabla_{\theta}f_{\hat{\theta}}(x),$ (21)
assuming $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist, and that ${\mathcal{L}_{\theta}}^{(x,\lambda)}$ is continuously differentiable w.r.t. $\theta$.
Proof.
To improve readability we use $(x,\lambda)$ in this proof to refer to fact that evaluation $x$ is passed to the regularizing term (i.e. always $f(x)$). Consider the derivative w.r.t. $\theta$ of the prediction-regularized objective at the optimum,
@@ Appendix A PROOFS
Theorem 4.2 The optimal regularization term defined by Eq. 13 is $\hat{r}=\frac{1}{M}\sum_{m=1}^{M}f(\hat{x}_{m})$ when $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist and ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
@@
Table 2: Notation Summary
Symbol Definition
$\theta$ Network parameters.
$\mathcal{L}_{\theta}$ Log joint distribution parameterized by $\theta$.
$\hat{\theta}$ Maximum a posteriori parameters under the joint distribution (Eq. 1).
$f_{\hat{\theta}}$ Network output with parameters $\hat{\theta}$.
$\nabla_{\theta}f_{\hat{\theta}}$ Derivative of network output w.r.t. $\theta$ evaluated at $\theta=\hat{\theta}$.
${\mathcal{L}_{\theta}}^{(t_{\theta},\lambda)}$ Log joint distribution parameterized by $\theta$ and regularized by $\lambda t_{\theta}$, e.g. $t_{\theta}$ could be a network output $f_{\theta}(x)$ (Eq. 8) or fine-tuning regularizer $r$ (Eq. 13).
$\hat{\theta}^{(t_{\theta},\lambda)}$ Parameters that maximize ${\mathcal{L}_{\theta}}^{(t_{\theta},\lambda)}$, e.g., see Eq. 9.
$\nabla_{\lambda}f_{\hat{\theta}^{(t_{\theta},\lambda)}}$ Derivative of network output under parameters $\hat{\theta}^{(t_{\theta},\lambda)}$ w.r.t. regularization weight $\lambda$. Note that the optimal network parameters themselves are a function of $\lambda$, making the overall term a non-linear expression involving $\nabla_{\lambda}f_{\hat{\theta}}$ (see Eq. 21).
Appendix B ALGORITHMS FOR REGULARIZATION VARIATION (RegVar)
In this section we expand on variations of the RegVar strategy and provide relevant algorithms where applicable.
B.1 Pointwise RegVar
The algorithm that follows most directly from Theorem 1 targets the uncertainty calculation for an arbitrary evaluation point $x$ (not necessarily from the training or test set) and is given in Algorithm 2. The computational burden of recalculating the optimal regularized parameters for every different evaluation input means that this approach is not scalable to large models, motivating extensions to the regularization term which we consider next.
Algorithm 2 Pointwise RegVar
  Warm start $\hat{\theta}^{(x,\lambda)}\leftarrow\hat{\theta}$
  Initialize step counter $j=0$
  repeat
     Sample training example $(x_{i},y_{i})\sim\mathcal{D}$
     Follow stochastic gradient $\hat{\theta}^{(x,\lambda)}\leftarrow\hat{\theta}^{(x,\lambda)}+N\gamma(j)\nabla_{\theta}\tilde{\mathcal{L}}_{\theta}^{(x,\lambda)}(x_{i},y_{i})$
     $j\leftarrow j+1$
  until $\hat{\theta}^{(x,\lambda)}$ converges
B.2 Data-Augmentation Objective
When the likelihood is Gaussian, the RegVar regularization term can be absorbed into the likelihood. In particular, completing the square in Eq. 16, by pulling the term $\frac{\lambda\sigma^{2}}{nM}(\frac{\lambda\sigma^{2}}{2nM}-y_{i})$ from a constant that does not depend on $\theta$, yields the objective,